OpenStudy (anonymous):

A right triangle has an angle that measures 34, and the adjacent side measures 17. What is the length of the hypotenuse to the nearest tenth?

OpenStudy (campbell_st):

 use cos34 = 17/x then x = 17/cos34 hint... the hypotenuse is the longest side so you answer needs to be longer than 17